The week of Oct 13

Our electrical tech began work on Virus enclosure this week. We will be installing all of wiring and breakers in the enclosure. One of our telescope operators is working with one of our programers to sort out the remaining problems with the wiring to the anemometers. The next step will be to set the zero point of the wind vein in the software which requires them to point the anemometers in a known direction and note the position read out. We also installed a new digital rain gauge this week. We will now be able to record when we have rain events in the weather archive. Work on the TCS continues and some bugs in the RA and DEC offsets were found and removed. Offsetting in X,Y, rho are still to be done.

This was supposed to be the last week for work on the Facility Thermal Management Project. Veliz Construction finished the back stairway past the K-hut to the liquid nitrogen slab. This was a requirement by the people who will be filling the liquid nitrogen tank. A fair amount of time was spent putting insulation around pipes, cleaning up the site and small projects. The consulting engineers from El Paso came out on Thursday to generate the punch list of remaining items for the contractors.

Leave a Reply

Your email address will not be published. Required fields are marked *